3.0cm, s^-1 + (9.8ms^-2 x 2.0ms)
Please can you help me with this as quick as possible and can the answer be in ms^-1 Thanks :)
1
Expert's answer
2015-09-08T01:27:21-0400
3.0cm, s^-1 + (9.8ms^-2 x 2.0ms) =
3.0*0.01 m/s + 9.8 m/s^2 * 2*m*s =
=0.03 m/s + 19.6 m^2/s